OpenStudy (anonymous):

Suppose the radius of a spherical balloon is growing at the rate of 50 cm/s, so how quickly is the volume increasing when the radius is 6 cm?

jigglypuff314 (jigglypuff314):

well from the previous question we know that V = (4/3) pi r^3 then derivative of V = V' = 4 pi r^2 r' we were then given that the radius increases at a rate of 50 so rate of radius = derivative of radius = r' = 50 and radius = 6 then plug that in so V' = 4 pi (6)^2 (50) = ?

OpenStudy (anonymous):

7200pi

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

Thanks, buddy!

jigglypuff314 (jigglypuff314):

you're welcome :)